detail for solution
Dr. Oz then pointed out the part that really caught my attention. They did CT scans of the chest in all the patients. All
Answer #1

Given that,
sample one, x1 =17, n1 =31, p1= x1/n1=0.55
sample two, x2 =25, n2 =31, p2= x2/n2=0.81
null, Ho: p1 = p2
alternate, H1: p1 != p2
level of significance, α = 0.05
from standard normal table, two tailed z α/2 =1.96
since our test is two-tailed
reject Ho, if zo < -1.96 OR if zo > 1.96
we use test statistic (z) = (p1-p2)/√(p^q^(1/n1+1/n2))
zo =(0.548-0.81)/sqrt((0.677*0.32(1/31+1/31))
zo =-2.17
| zo | =2.17
critical value
the value of |z α| at los 0.05% is 1.96
we got |zo| =2.173 & | z α | =1.96
make decision
hence value of | zo | > | z α| and here we reject Ho
p-value: two tailed ( double the one tail ) - Ha : ( p != -2.1734 ) = 0.0297
hence value of p0.05 > 0.0297,here we reject Ho
ANSWERS
---------------
1.
null, Ho: p1 = p2
alternate, H1: p1 != p2
test statistic: -2.17
critical value: -1.96 , 1.96
decision: reject Ho
p-value: 0.0297
2.
yes,
statistically significant
3.
we have enough evidence to support the claim that difference in proportion to show a lot of people whether this is the right way or the wrong way to go.